Several years after reengineering its production​ process, Biltmore Corporation hired a new​ controller, Georgia Taylor. She developed an ABC system very similar to the one used by Biltmore​'s chief rival. Part of the reason Taylor developed the ABC system was because Biltmore​'s profits had been declining even though the company had shifted its product mix toward the product that had appeared most profitable under the old system. Before adopting the new ABC​ system, the company had used a plantwide overhead rate based on direct labor hours that was developed years ago.

For the upcoming​ year, Biltmore​'s budgeted ABC manufacturing overhead allocation rates are as​ follows:

Activity Allocation Base Activity Cost Allocation Rate
Materials Handling Number of parts $4.00 per part
Machine Setup Number of setups $450.00 per setup
Insertion of parts Number of parts $26.00 per part
Finishing Finishing direct labor hours $50.00 per hour

The number of parts is now a feasible allocation base because Biltmore recently installed a plantwide computer system. Biltmore produces two wheel​ models: Standard and Deluxe. Budgeted data for the upcoming year are as​ follows:

Standard Deluxe
Parts per wheel 4.0 6.0
Setups per 1,000 wheels 20.0 20.0
Finishing direct labor hours per wheel 1.1 2.4
Total direct labor hours per wheel 2.6 3.2

The​ company's managers expect to produce 1,000 units of each model during the year.

1.

Compute the total budgeted manufacturing overhead cost for the upcoming year. Enter rate to two decimal places.

2.

Compute the manufacturing overhead cost per wheel of each model using ABC.

3.

Compute the​ company's traditional plantwide overhead rate. Use this rate to determine the manufacturing overhead cost per wheel under the traditional system.
Answer #1

1.Total budgeted manufacturing overhead cost for the upcoming year:

Activity allocation base Budgeted quantity of cost allocation base allocation rate Budgeted indirect cost
Materials handling Number of parts (4*1000) +(6*1000) = 10,000 $4 $40,000
Machine setups Number of setups (20) +(20) =>40 $450 $18,000
Insertion of parts Number of parts (4*1000) +(6*1000) =>10,000 $26 $260,000
Finishing Finishing direct labor hours (1.1*1000) +(2.4*1000)=>3500 $50 $175,000
total budgeted indirect cost $493,000

Requirement 2:

manufacturing overhead cost per wheel of each model:

Activity cost allocation rate quantity of cost allocation base used Allocated activity cost per wheel
Standard Deluxe standard Deluxe
Materials handling $4 per part 4 6 $16 $24
Machine setup $450 per set up (20 setups /1000)=0.02 (20 setups/1000)=0.02 $9 $9
Insertion of parts $26 per part 4 6 $104 $156
Finishing $50 per hour 1.1 2.4 $55 $120
Total ABC allocated indirect cost $184 $309

3. Traditional plant-wide overhead rate:

total direct labour hours = (2.6*1000) +(3.2*1000) =>5800 hours

(total overhead ) /(total labor hours) = Plantwide overhead rate

($493,000) /(5800 hours) =>$85.per direct labor hour

NOw, manufacturing overhead cost per wheel using a plant-wide overhead rate:

..................................Direct labor hours * .................plant-wide rate.=.. manufacturing overhead:

Standard 2.6....................X $85 $221
Deluxe 3.2....................X $85 $272